OpenStudy (anonymous):

how to solve Rtot from 1/R+1/3R=1/Rtot

OpenStudy (lastdaywork):

Which grade do you study in ?

OpenStudy (anonymous):

6

OpenStudy (lastdaywork):

So you have not been taught to add fractions ?

OpenStudy (anonymous):

No not yet...

OpenStudy (anonymous):

I want to learn new things

OpenStudy (lastdaywork):

Its simple; to add \[\frac{ a }{b }+\frac{ c }{ d }\] first make their denominator same. This can be done by finding LCM (Least Common Multiple) of the denominators (namely b and d) Or you can simply make each denominator equal to the product of the two denominators (i.e. b*d) Here, I will use the second approach. A fraction remains same when it is multiplied and divided by the same number; so \[\frac{ a }{b }+\frac{ c }{ d }=\frac{ a*d }{ b*d }+\frac{ c*b }{ d*b }\] When the denominators are same, you can add the (new) numerators like integers. So \[\frac{ a }{b }+\frac{ c }{ d }=\frac{ a*d+c*b }{ b*d }\] As an exercise - solve \[\frac{ 1 }{ R }+\frac{ 1 }{ 3R }\]

OpenStudy (lastdaywork):

@AlexKrakoq You wanted to learn how to do the calculation or just the answer ??